1. A job with a 90% learning curve required 20 hours for the initial unit. The eighth unit should require approximately how many hours?


TN = T1C


2. A job requires 20 hours for the initial unit and 12.8 hours for the fourth unit. What is the learning rate?


TN = T1C


3. A job with a 95% learning curve required 19 hours for the second unit. The first unit should require approximately how many hours?


TN = T1C




1
Expert's answer
2021-01-24T18:43:49-0500

"T_1\times L^n=T_n," time required for the nth unit

where "T_1=" unit time of the first unit

"L=" learning curve raten

"n=" number of times T is doubled


1.

Hours required for unit 8 "=20\times(0.9)^3=14.58" hours.


2.

Hours required for unit 4 "=20\times(L)^2=12.8" hours.


"L^2=0.64"

"L=0.8"

The learning rate is "80\%."


3.

Hours required for unit 2 "=T_1\times(0.95)^1=19" hours.


"T_1=\dfrac{19}{0.95}=20(hours)"

A job with a 95% learning curve required 20 hours for the initial unit.